OpenStudy (anonymous):

can someone solve this exponential equation 3^x+1=9^2x-1

OpenStudy (lgbasallote):

first..you need to know that 9 = 3^2 so 9^2x-1 = 3^2(2x-1) do you have an idea on what to do now?

OpenStudy (anonymous):

once the base is the same, then the exponents would drop to solve 2(2x-1)=x+1

OpenStudy (lgbasallote):

yes. very good! :DDD

OpenStudy (anonymous):

so I get 4x-2=x+1. Solve to get x=1

OpenStudy (lgbasallote):

correct.

OpenStudy (anonymous):

I have 2 more I need help with

OpenStudy (lgbasallote):

go ahead

OpenStudy (anonymous):

8^x^2-2x=.5

OpenStudy (anonymous):

Thanks Luis. How about this one 8^x^2-2x=.5

OpenStudy (anonymous):

one moment? Great. I've been working for almost 20 minutes

OpenStudy (anonymous):

no. x^2-2x. 2x is part of the exponent to the 8. The 2 is the exponent to the x. Does that make any sense?

OpenStudy (anonymous):

WOW. Very impressive. I have no idea what it means. However, the number after the equal sign is .5 or 1/2 not 5.

OpenStudy (anonymous):

This is my first time here. What does the medal mean?

OpenStudy (anonymous):

Thanks! Here is my last one if you don't mind: (2^x^2+1)/(2^x-1)=16